# latex‘!h’ float specifier changed to ‘!ht’

### v1

@@ Abstract
Graph Convolutional Networks (GCNs) have become a pivotal method in machine learning for modeling functions over graphs. Despite their widespread success across various applications, their statistical properties (e.g. consistency, convergence rates) remain ill-characterized. To begin addressing this knowledge gap, in this paper, we provide a formal analysis of the impact of convolution operators on regression tasks over homophilic networks. Focusing on estimators based solely on neighborhood aggregation, we examine how two common convolutions—the original GCN and GraphSage convolutions—affect the learning error as a function of the neighborhood topology and the number of convolutional layers. We explicitly characterize the bias-variance trade-off incurred by GCNs as a function of the neighborhood size and identify specific graph topologies where convolution operators are less effective. Our theoretical findings are corroborated by synthetic experiments, and provide a start to a deeper quantitative understanding of convolutional effects in GCNs for offering rigorous guidelines for practitioners.
@@ 1 Introduction
Graph Convolutional Networks (GCNs [10, 17]) have become one of the preferred tools for modeling, analyzing, and predicting signals on graphs. Despite their impressive success on academic benchmarks, several fundamental issues limit their broader applicability and reliability in real-world scenarios. In particular, while graphs are highly diverse in their properties (e.g., sparsity, degree distribution, node feature types) and the relationships they encode, GCNs are often suggested as a one-size-fits-all approach. This leaves practitioners with the challenge of selecting an appropriate GCN convolution and architecture for their specific task. Without a clear understanding of the inductive biases these convolutions encode and their expected performance in relation to dataset properties, choosing among the 66 available convolution options in pytorch geometric [25] can become a daunting task. To begin bridging this knowledge gap, this paper investigates the properties of graph convolutions, and relates it to graph or neighborhood properties, focusing specifically on regression tasks over networks.
@@
In this paper, we consider the problem of network regression under fixed design. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ denote a graph with vertex set $\mathcal{V}$ and edge set $\mathcal{E}$. Assume that at each node $i\in\{1,\ldots,n\}$, we observe a corresponding real-valued response variable $Y_{i}$, which is generated according to the following mechanism:
@@
where the noise variables ${\varepsilon}_{i}$ are assumed to be uncorrelated, centered, and have variance equal to one. Let $\mathbf{Y}:=(Y_{1},\ldots,Y_{n})^{\top}\in\mathbb{R}^{n}$ denote the vector of responses from all nodes in the graph $\mathcal{G}$, ${\bm{{\varepsilon}}}:=({\varepsilon}_{1},\ldots,{\varepsilon}_{n})^{\top}$ the noise vector, and $\mathbf{f^{*}:}=(f_{1}^{*},\ldots,f_{n}^{*})^{\top}$ the regression vector. Our goal is to recover $\mathbf{f^{*}}$ from the observations $\mathbf{Y}$. We also assume that $\mathcal{G}$ is a homopholic graph, so that neighboring nodes are expected to have similar signal: $f_{i}^{*}\approx f_{j}^{*}$, if $(i,j)\in\mathcal{E}.$
@@
with ${\mathcal{F}}$ a prespecified function class. To incorporate information about the graph, graph-based regularization methods consider instead a penalized objective of the form:
$\mathbf{\hat{f}}_{\lambda}$ $:=\mathop{\rm argmin}_{g}\big{\|}\mathbf{Y}-g(\mathbf{Y})\big{\|}_{2}^{2}+\lambda\sum_{(i,j)\in{\mathcal{E}}}\text{Pen}\big{(}g(\mathbf{Y})_{i}-g(\mathbf{Y})_{j}\big{)},$
@@
where Pen denotes a function penalizing differences along edges of the graph.
@@
When the graph is endowed with node features $X\in\mathbb{R}^{n\times p}$, and the estimator $\mathbf{\hat{f}}$ is chosen to be a function of these features (e.g. $\mathbf{\hat{f}}=\mathbf{\hat{f}}(X)$), the task is usually known as node regression. When the estimator is constructed using information solely from the observed values of $\mathbf{Y}$, the problem boils down to a denoising task usually known as graph-trend filtering [15, 35], an extension of the classical Gaussian Sequence model [26, 36] to accommodate graph structure. In this paper, we aim to study the performance of graph convolutionals to provide reliable estimates for $\mathbf{f^{*}}$ in the denoising setting.
Notations.
@@
Let $|\mathcal{A}|$ represent the cardinality of a set $\mathcal{A}$. We denote a graph with vertex set $\mathcal{V}$ and edge set $\mathcal{E}$ by $\mathcal{G}=(\mathcal{V},\mathcal{E})$. Without loss of generality, we assume that the vertices are indexed from 1 to $n$, so $\mathcal{V}=\{1,\ldots,n\}$ and $|\mathcal{V}|=n$. Let $A=(A_{ij})_{i,j\in\mathcal{V}}\in{\mathbb{R}}^{n\times n}$ denote the binary adjacency matrix, which encodes the edge structure of the graph $\mathcal{G}$ as follows:
@@
For any $k=1,2,\ldots$, we further denote by $\mathcal{N}^{k}(i)$ all nodes that are at most $k$ hops away from node $i$. We denote the neighborhood of node $i$ by $\mathcal{N}(i)$. Note that $\mathcal{N}^{1}(i)=\mathcal{N}(i)\cup\{i\}$. For any $i\in\mathcal{V}$, let $d_{i}$ denote the edge degree of node $i$, specifically $d_{i}=|\mathcal{N}(i)|$. The degree matrix $D$ is the $n\times n$ diagonal matrix with diagonal entries
$D_{ii}=d_{i},\quad i=1,\ldots,n.$
The graph Laplacian is denoted by
$\mathbf{L}=D-A,$
@@
and is a positive semi-definite and symmetric matrix [4]. For a vector ${\bf{v}}=(v_{1},\ldots,v_{n})\in{\mathbb{R}}^{n}$, $\|{\bf{v}}\|_{2}^{2}=\sum_{i=1}^{n}v_{i}^{2}$ denotes the squared Euclidean norm and the notation $({\bf{v}})_{i}=v_{i}$ represents the $i$-th component of the vector. For a square matrix $M$, let $\|M\|_{F}$ denote the Frobenius norm and $\mathop{\rm Tr}\nolimits(M)$ denote the trace.
Related literature.
The theory for node regression has been derived from different viewpoints. Asymptotic viewpoint assumes that the number of vertices $n=|\mathcal{V}|$ tends to infinity, so that the Laplacian matrix is used to approximate the Laplace-Beltrami operator on an underlying manifold [1, 19, 29] on which the function $\mathbf{f^{*}}$ is assumed to be smooth [11, 34]. Instead of assuming a growing graph, one can alternatively work with a fixed graph where, for a randomly selected subset of nodes, the response vectors are masked/unobserved [17, 6]. A typical assumption consists in assuming that neighboring nodes will have the same response – in other words, assuming that the graph is homophilic. In this setting, several strategies leveraging the graph structure to guide the inference can be deployed: (a) Laplacian regularization, (b) $\ell_{1}$ regularization and (c) graph convolutions.
Laplacian regularization (also known as Laplacian smoothing [30], a special instance of Tikhonov regularisation) is perhaps one of the better studied version of the problem. As described in a previous paragraph, these methods usually consider a smooth function (where the smoothness of the function $\mathbf{f^{*}}$ is consequently defined via the eigenvalue decay of the graph Laplacian), and use the convergence of the graph Laplacian to the Laplace-Beltrami operator to establish the consistency and convergence rates of the proposed estimator [11, 34]. Several research groups have explored similar concepts. [2] provides bounds on the generalization error for Tikhonov regularization and interpolated regularization. Work by [18] derives posterior contraction rates for $\mathbf{f^{*}}$ using a Bayesian formalism. The rates depend on the eigenvalue decay of the graph Laplacian and a smoothness index that is defined via a Sobolev-type space based on the graph Laplacian. More recently, work by Green and coauthors [8, 9] derives optimal convergence rates for Laplacian-based regularization in the fixed design setting.
Graph-trend filtering (also known as “graph total-variation” or “fused lasso”) estimators use an $\ell_{1}$ penalty to guide the inference, so that estimators are derived as the solution of the following optimisation problem:
$\mathbf{\hat{f}}\in\text{argmin}_{\mathbf{x}\in\mathbb{R}^{n}}\|\mathbf{Y}-\mathbf{x}\|_{2}^{2}+\lambda\sum_{i\sim j}|x_{i}-x_{j}|.$
Recent work, see [15, 24] has characterized the convergence rate of these estimators as a function of the underlying graph smoothness, defined this time in the amount of variation registered along the graph edges: $\sum_{i\sim j}|f^{*}_{i}-f^{*}_{j}|.$ While this approach is attractive for approximative near piece-wise constant functions over graphs, estimators are often more computational intensive than Laplacian based estimators.
Finally, Graph Convolution Neural Networks [17] are a more recent addition to the set of methods. While typically used for classification, this method is increasingly used for node regression tasks. The theoretical understanding of GCNs models remains quite limited. For example, [32] analyses the stability of single-layer GCNs models and derives their generalization guarantees, showing that, in regression problems, the generalization gap between training error and testing error decreases at a sub-linear rate as the number of training samples increases. The work in [12, 5] demonstrates that GCNs can be interpreted as a simplification of spectral-type GNNs that utilize the graph Laplacian. In [20], it is shown that graph convolution is essentially a form of Laplacian smoothing. [33] investigates the classification performance of graph neural networks with graph-polynomial features under a general contextual stochastic block model.
Contrary to earlier work, in this paper we aim to provide statistical insight into the regression setting. We analyze a widely used class of GNNs, Graph Convolutional Networks (GCNs), which we introduce in depth in the next section.
2 Graph convolutional networks (GCNs)
@@
While graph neural networks (GNNs) is a broad term encompassing various network structures for graph learning, graph convolutional networks (GCNs) [17] form a specific class of GNNs. GCNs take the response vector $\mathbf{Y}$ and the adjacency matrix $A$ as the input. Let $L$ be the number of GCN layers. For each $\ell\in\{0,\ldots,L-1\}$, the output of the layers can be defined recursively via
@@
Here, $\sigma$ denotes componentwise application of the ReLU activation function, $W\in\mathbb{R}$ is a weight parameter, and
@@
for the adjacency/degree matrix with added self-loops $\tilde{A}$ and $\tilde{D},$ that is, $\tilde{A}=A+I_{n}$ and $\tilde{D}$ is the diagonal $n\times n$ matrix with diagonal entries $\tilde{D}_{ii}=\sum_{j}\tilde{A}_{ij}=d_{i}+1.$
The definition can be extended to the case where at each node we do not only observe one number but a feature vector of length $p$. In this case, $\mathbf{Y}$ becomes a $n\times p$ matrix and $W$ can be chosen as $p\times p$ matrix.
The formulation of (2) originates from [17]. Its motivation is to provide a first-order approximation for the localized spectral filters introduced in [5] for signal processing, suggesting that GCNs should be applied in the fixed design framework.
In addition to the widely used propagation operator $T$, this paper also discusses the original normalization approach, that replaces $T$ by
@@
with $P=T$, or $P=S$, and $W_{0},\ldots,W_{L-1}\in\mathbb{R}$ are the weight parameters in each layer. Empirical studies demonstrate that, compared to nonlinear activation functions (2), linear GCNs as defined in (5) do not have a detrimental effect on accuracy in many downstream applications [37]. In [23], the authors provide theoretical evidence that non-linearity does not enhance the expressive power of GCNs. The linear setting has also been considered in [33] and referred to as poly-GNN. Further discussion comparing linear GCNs to the original GCNs can be found in [21, 22].
@@
with some $W\in\mathbb{R}$. We denote the class of functions described in (6) as $\mathcal{F}_{\text{GCN}}(L)$ and base our analysis on this class.
@@
Several studies have begun investigating the capabilities and limitations of Graph Neural Networks (GNNs). Most analyses focus on comparing different classes of GNN architectures using the Weisfeiler-Lemann test to distinguish graph structures [38] or on performance in classification tasks [7]. While issues such as oversmoothing in deeper GNNs have been explored [20, 23], there is a noticeable gap in evaluating the consistency of GNN-based estimators. In particular, little has been done to analyze the bias-variance trade-off concerning neighborhood size and topology. This gap highlights the need for a deeper understanding of how graph convolution operations affect the statistical properties of GNNs, motivating our investigation.

### v2

# Understanding the Effect of GCN Convolutions in Regression Tasks

@@ Abstract
Graph Convolutional Networks (GCNs) have become a pivotal method in machine learning for modeling functions over graphs. Despite their widespread success across various applications, their statistical properties (e.g., consistency, convergence rates) remain ill-characterized. To begin addressing this knowledge gap, we consider networks for which the graph structure implies that neighboring nodes exhibit similar signals and provide statistical theory for the impact of convolution operators. Focusing on estimators based solely on neighborhood aggregation, we examine how two common convolutions—the original GCN and GraphSAGE convolutions—affect the learning error as a function of the neighborhood topology and the number of convolutional layers. We explicitly characterize the bias-variance type trade-off incurred by GCNs as a function of the neighborhood size and identify specific graph topologies where convolution operators are less effective. Our theoretical findings are corroborated by synthetic experiments, and provide a start to a deeper quantitative understanding of convolutional effects in GCNs for offering rigorous guidelines for practitioners.
@@ 1 INTRODUCTION
Graph Convolutional Networks (GCNs) have become one of the preferred tools for modeling, analyzing, and predicting signals on graphs (Hamilton et al.,, 2017; Kipf and Welling,, 2017). Despite their impressive success on academic benchmarks, several fundamental issues limit their broader applicability and reliability in real-world scenarios. In particular, while graphs are highly diverse in their properties (e.g., sparsity, degree distribution, node feature types) and the relationships they encode, GCNs are often suggested as a one-size-fits-all approach. This leaves practitioners with the challenge of selecting an appropriate convolution and architecture for their specific task. Without a clear understanding of the inductive biases these convolutions encode and their expected performance in relation to dataset properties, choosing among the sixty-six available convolution options in pytorch geometric can become a daunting task (Paszke et al.,, 2019). To bridge this knowledge gap, this paper examines the properties of graph convolutions and their relationship to the graph’s neighborhood characteristics, with a particular focus on regression tasks over networks.
@@
We consider the problem of network regression under fixed design. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ denote a graph with vertex set $\mathcal{V}$ and edge set $\mathcal{E}$. Without loss of generality, we index the vertices from 1 to $n$, so $\mathcal{V}=\{1,\ldots,n\}$ and $|\mathcal{V}|=n$. Assume that at each node $i\in\{1,\ldots,n\}$, we observe a corresponding real-valued response variable $Y_{i}$, which is generated according to the following mechanism:
@@
where the noise variables ${\varepsilon}_{i}$ are assumed to be uncorrelated, centered, and have variance equal to one. Let $\mathbf{Y}:=(Y_{1},\ldots,Y_{n})^{\top}\in\mathbb{R}^{n}$ denote the vector of responses from all nodes in the graph $\mathcal{G}$, ${\bm{{\varepsilon}}}:=({\varepsilon}_{1},\ldots,{\varepsilon}_{n})^{\top}\in\mathbb{R}^{n}$ the noise vector, and ${\mathbf{f^{*}:}=(f_{1}^{*},\ldots,f_{n}^{*})^{\top}\in\mathbb{R}^{n}}$ the regression vector, which we also refer to as the signals. Throughout the paper, we assume that neighboring nodes in $\mathcal{G}$ admit similar signals: $f_{i}^{*}\approx f_{j}^{*}$, if $(i,j)\in\mathcal{E}$. Our goal is to recover $\mathbf{f^{*}}$ from the observations $\mathbf{Y}$.
@@
with ${\mathcal{F}}$ being a prespecified function class. To incorporate information about the graph, graph-based regularization methods consider instead a penalized objective of the form:
$\mathbf{\hat{f}}_{\lambda}:=$ $\mathop{\rm argmin}_{g\in{\mathcal{F}}}\big{\|}\mathbf{Y}-g(\mathbf{Y})\big{\|}_{2}^{2}$
$+\lambda\sum_{(i,j)\in{\mathcal{E}}}\text{Pen}\big{(}g(\mathbf{Y})_{i}-g(\mathbf{Y})_{j}\big{)}$
@@
where Pen denotes a function penalizing differences along edges of the graph, and $g(\mathbf{Y})_{i}$ represents the $i$-th component of $g(\mathbf{Y})\in\mathbb{R}^{n}$.
@@
When the graph is endowed with a node feature matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$, and the estimator $\mathbf{\hat{f}}$ is chosen to be a function of these features (e.g. $\mathbf{\hat{f}}=\mathbf{\hat{f}}(\mathbf{X})$), the task is usually known as node regression. When the estimator is constructed using information solely from the observed values of $\mathbf{Y}\in\mathbb{R}^{n}$, the problem boils down to a denoising task usually known as graph-trend filtering (Hütter and Rigollet,, 2016; Wang et al.,, 2016), an extension of the classical Gaussian sequence model (Rigollet,, 2015; Wasserman,, 2006) to accommodate graph structure. In this paper, we aim to study the performance of graph convolutions in providing reliable estimates for $\mathbf{f^{*}}$ under the denoising setting. Specifically, the feature $X_{i}$ associated with each node $i$ is simply given by $X_{i}=i$.
Contributions.
The main contributions of this paper are summarized as follows:
  • We establish a bias-variance type trade-off for the resulting denoising estimator in terms of the number of convolutional layers. This trade-off highlights how increasing the number of layers can reduce variance while potentially introducing greater bias, thereby affecting the overall estimation accuracy.
  • By showing that the variance of the GCN is a weighted sum over paths, we relate the variance of GCNs to that of the uniformly local averaging estimator and introduce the novel walk analysis approach to investigate the problem. This method may pave the way for further statistical studies of GCNs.
  • Based on the proposed walk analysis, we further quantitatively demonstrate the different variance decay behaviors under distinct local topologies of the graph. Specifically, we show that when high edge degree nodes are connected to low edge degree nodes this can dominate the variance and result in much slower overall variance decay.
  • Additionally, we provide a concrete example of the graph structure to illustrate the over-smoothing phenomenon, which serves as a prototype to explain why simply stacking layers leads to poor performance in practice. Whereas previous work has provided evidence of this phenomenon for asymptotically large graphs and classification settings (Oono and Suzuki,, 2020; Cai and Wang,, 2020), we identified a non-asymptotic regression setting.
2 RELATED WORK
The theory for node regression has been derived from different viewpoints. Asymptotic viewpoint assumes that the number of vertices $n$ tends to infinity, so that the Laplacian matrix is used to approximate the Laplace-Beltrami operator on an underlying manifold (Belkin and Niyogi,, 2003; Lafon,, 2004; Singer,, 2006) on which the function $\mathbf{f^{*}}$ is assumed to be smooth (Hein,, 2006; Von Luxburg et al.,, 2008). Instead of assuming that the number of vertices $n$ tends to infinity, one can alternatively work with a fixed graph where, for a randomly selected subset of nodes, the response vectors are masked/unobserved (Kipf and Welling,, 2017; Donnat et al.,, 2024). A typical assumption consists in assuming that neighboring nodes will have the same response. In this setting, several strategies leveraging the graph structure to guide the inference can be deployed: (a) Laplacian regularization, (b) $\ell_{1}$ regularization and (c) graph convolutions.
Laplacian regularization (also known as Laplacian smoothing (Smola and Kondor,, 2003), a special instance of Tikhonov regularization) is perhaps one of the better studied version of the problem. As described in a previous paragraph, these methods usually consider a smooth function (where the smoothness of the function $\mathbf{f^{*}}$ is consequently defined via the eigenvalue decay of the graph Laplacian), and use the convergence of the graph Laplacian to the Laplace-Beltrami operator to establish the consistency and convergence rates of the proposed estimator (Hein,, 2006; Von Luxburg et al.,, 2008). Several research groups have explored similar concepts. Belkin et al., (2004) provides bounds on the generalization error for Tikhonov regularization and interpolated regularization. Work by Kirichenko and van Zanten, (2017) derives posterior contraction rates for $\mathbf{f^{*}}$ using a Bayesian formalism. The rates depend on the eigenvalue decay of the graph Laplacian and a smoothness index that is defined via a Sobolev-type space based on the graph Laplacian. More recently, work by Green and coauthors (Green et al.,, 2021, 2023) derives optimal convergence rates for Laplacian-based regularization in the fixed design setting.
Graph-trend filtering (also known as “graph total-variation” (Hütter and Rigollet,, 2016) or “fused lasso” (Tibshirani et al.,, 2005; Tibshirani and Taylor,, 2011)) estimators use an $\ell_{1}$ penalty to guide the inference, so that estimators are derived as the solution of the following optimization problem:
$\mathbf{\hat{f}}\in\text{argmin}_{\mathbf{x}\in\mathbb{R}^{n}}\|\mathbf{Y}-\mathbf{x}\|_{2}^{2}+\lambda\sum_{(i,j)\in\mathcal{E}}|x_{i}-x_{j}|.$
Recent work (e.g., see Donnat et al., (2024); Hütter and Rigollet, (2016); Padilla et al., (2018)) has characterized the convergence rate of these estimators as a function of the underlying graph smoothness, defined by the amount of variation along the graph edges: $\sum_{(i,j)\in\mathcal{E}}|f^{*}_{i}-f^{*}_{j}|.$ While this approach is attractive for approximating near piecewise-constant functions over graphs, the estimators are often more computationally intensive than Laplacian-based estimators.
Finally, Graph convolutional networks (Kipf and Welling,, 2017) are a more recent addition to the set of methods. While typically used for classification, this method is increasingly used for node regression tasks. The theoretical understanding of GCN models remains quite limited. For example, Verma and Zhang, (2019) analyzes the stability of single-layer GCN models and derives their generalization guarantees, showing that, in regression problems, the generalization gap between training and testing errors decreases at a sub-linear rate as the number of training samples increases. The work in Henaff et al., (2015) and Defferrard et al., (2016) demonstrates that GCNs can be interpreted as a simplification of spectral-type Graph Neural Networks (GNNs) that makes use of the graph Laplacian. In Li et al., (2018), it is shown that graph convolution is essentially a form of Laplacian smoothing. Vinas and Amini, (2024) investigates the classification performance of GNNs with graph-polynomial features under a general contextual stochastic block model.
Contrary to earlier work, in this paper we aim to provide statistical insight into the regression setting. We analyze a widely used class of GNNs, namely GCNs, which we discuss in detail in the next section.
Notation.
@@ 2 RELATED WORK
Let $A=(A_{ij})_{i,j\in\mathcal{V}}\in{\mathbb{R}}^{n\times n}$ denote the binary adjacency matrix, which encodes the edge structure of the graph $\mathcal{G}$ as follows:
@@
For any $k=1,2,\ldots$, let $\mathcal{N}^{k}(i)$ represent the nodes within $k$ hops of node $i$. The neighborhood of node $i$ is denoted by $\mathcal{N}(i)$, with $\mathcal{N}^{1}(i)=\mathcal{N}(i)\cup\{i\}$. For $i\in\mathcal{V}$, let $d_{i}:=|\mathcal{N}(i)|$ be the edge degree of node $i$. The degree matrix $D$ is the $n\times n$ diagonal matrix with diagonal entries $D_{ii}=d_{i},$ $i=1,\ldots,n.$ For a vector ${\bf{v}}=(v_{1},\ldots,v_{n})\in{\mathbb{R}}^{n}$, $\|{\bf{v}}\|_{2}^{2}=\sum_{i=1}^{n}v_{i}^{2}$ denotes the squared Euclidean norm and the notation $({\bf{v}})_{i}=v_{i}$ represents the $i$-th component of the vector. For a square matrix $M$, let $\|M\|_{F}$ denote the Frobenius norm and $\mathop{\rm Tr}\nolimits(M)$ denote the trace.
3 GRAPH CONVOLUTIONAL NETWORKS
@@
While GNNs is a broad term encompassing various network structures for graph learning, GCNs (Kipf and Welling,, 2017) form a specific class of GNNs. GCNs take the response vector $\mathbf{Y}\in{\mathbb{R}}^{n}$ and the adjacency matrix $A\in{\mathbb{R}}^{n\times n}$ as the input. Let $L$ be the number of GCN layers. For each $\ell\in\{0,\ldots,L-1\}$, the output of the layers can be defined recursively via
@@
Here, $\sigma$ denotes the component-wise application of the ReLU activation function, $\mathbf{y}\in{\mathbb{R}}^{n}$ represents a realization of the random vector $\mathbf{Y}$, and $W\in\mathbb{R}$ is a weight parameter. The convolutional operator $T$ is defined as
@@
where $\tilde{A}:=A+I_{n}$ is the adjacency matrix with added self-loops, and $\tilde{D}$ is the diagonal $n\times n$ matrix with entries $\tilde{D}_{ii}=\sum_{j}\tilde{A}_{ij}=d_{i}+1.$
The above definition can be extended to the case where, at each node, we observe not just one number but a feature vector of length $p$. In this case, $\mathbf{Y}$ becomes an $n\times p$ matrix, and $W$ is a $p\times p$ matrix. For simplicity, we focus on the $p=1$ case for the rest of the paper.
The formulation of (2) originates from Kipf and Welling, (2017). Its motivation is to provide a first-order approximation for the localized spectral filters introduced in Defferrard et al., (2016) for signal processing, suggesting that GCNs should be applied in the fixed design framework.
In addition to the widely used propagation operator $T$, this paper also discusses an alternative normalization approach (known as the GraphSAGE convolution (Hamilton et al.,, 2017)), that replaces $T$ by
@@
with $P=T$, or $P=S$, and $W_{0},\ldots,W_{L-1}\in\mathbb{R}$ are the weight parameters in each layer. Empirical studies demonstrate that, compared to nonlinear activation functions (2), linear GCNs as defined in (5) do not have a detrimental effect on accuracy in many downstream applications (Wu et al.,, 2019). In Oono and Suzuki, (2020), the authors provide theoretical evidence that non-linearity does not enhance the expressive power of GCNs. Further discussion comparing linear GCNs to the original GCNs can be found in NT and Maehara, (2019) and NT et al., (2021).
@@
with some $W\in\mathbb{R}$. We denote the class of functions described in (6) as $\mathcal{F}_{\text{GCN}}(L)$ and base our analysis later on this class.
@@
Several studies have begun investigating the capabilities and limitations of Graph Neural Networks (GNNs), an umbrella term that includes Graph Convolutional Networks (GCNs) as an important example. Most analyses focus on comparing different classes of GNN architectures by using the Weisfeiler-Lemann test to distinguish between graph structures (Xu et al.,, 2018) or on performance in classification tasks (Garg et al.,, 2020). While issues such as over-smoothing in deeper GNNs have been explored (Li et al.,, 2018; Oono and Suzuki,, 2020), there is a noticeable gap in evaluating the consistency of GNN-based estimators. In particular, little has been done to analyze the bias-variance trade-off concerning neighborhood size and topology. This gap highlights the need for a deeper understanding of how graph convolution operations affect the statistical properties of GCNs, motivating our investigation.